# TPT-Bench: A Large-Scale, Long-Term and Robot-Egocentric Dataset for Benchmarking Target Person Tracking

### v1

@@ III-A1 Platform
To facilitate the collection of multi-sensor and egocentric data under robot kinematic constraints for non-robotics experts, we have developed a data collection platform illustrated in Fig. 2. This platform consists of a pushing cart equipped with a panoramic camera, a stereo camera, and a 3D LiDAR, operating under differential wheel constraints. The sensors are mounted at typical heights for mobile robots: the panoramic camera is placed approximately 73 cm above the ground, the 3D LiDAR at about 55 cm, and the stereo camera at approximately 45 cm with an elevation angle of around 30°. The extrinsic parameters of these sensors are well-calibrated and provided. We collected multi-sensor data, with detailed information provided in the supplementary materials.
Although our dataset includes multi-sensor data, we have annotated only the images from the panoramic camera to assess TPT performance using vision-based trackers. We will release all data in ROSBAG format, enabling future research on multi-modal TPT combining visual and depth modalities, natural following through imitation learning with multiple modalities and odometry, and more. Our data collection approach mirrors a human follower’s intention without delay, and all sensors are mounted similarly to the locations of most mobile robots. As such, the collected data can be directly used to train end-to-end networks for performing socially aware companionship, as demonstrated by recent advancements in learning-based social navigation  [29, 30].
Figure 2: Sensor suite and some third-view images. The dataset is demonstrated by a human pushing a sensor-equipped pushcart and following a target person in crowded environments indoors and outdoors. Blue represents the follower, and red represents the target person.
III-A2 Scenarios
To capture these long-term tracking challenges, we conducted person-following tasks across various indoor and outdoor locations during peak times, such as after classes, during lunch, and at dinner hours. The environments included canteens, laboratories, offices, university building halls, corridors, classrooms, supermarkets, roadside markets, crosswalks, sidewalks, metro stations, and more. Nine students (three females and six males) were recruited as target humans, navigating through these crowded environments and encountering issues like motion blur, varying lighting conditions, changing scales, and alterations in pose and clothing. These students also took on the role of followers, operating the sensor-equipped cart while adhering to social norms, ensuring they did not interrupt the leader or other pedestrians. Additionally, we occasionally carried out person-search tasks [31], which are essential for HRI, by asking the follower to fulfill the leader’s requirements first and then searching for the leader to resume the person-following behavior.
As a result, we compiled a large-scale, long-term, multi-modal, and diverse egocentric dataset consisting of 35 sequences captured in different locations, with multiple recordings taken at various times (e.g., daytime, sunset, and night). We then annotate the bounding boxes of the target person in panoramic images based on the standards of LaSOT [14], with a detailed annotation procedure and metadata involved in the supplementary materials.
Figure 3: Histogram of dataset statistics. In the “disappeared intervals” subfigure, durations over 30 seconds are excluded since they are considered part of the search procedure. The blue line represents the median.
III-B Data Statistics
In this benchmark, we aim to evaluate the challenges faced by TPT in long-term tracking across diverse scenarios. The relevant statistics are presented in Table I, with most metrics sourced from [16]. Additionally, we provide the distributions of tracking length, disappearance length, disappearance interval, number of disappearances, the number of surrounding people, target bounding box areas, bounding box variance within a sequence, and the distance of 3D bounding boxes from the sensor location, as shown in Fig. 3.
Compared to other person-tracking datasets (presented in Table I), our dataset provides significantly longer tracking sequences with an average sequence length of 394.3$\pm$138.4 seconds. These long-duration recordings capture numerous instances of target disappearance, with an average disappearance duration of 83.5 seconds per sequence. In some cases, the target may remain absent for more than 200 seconds, reflecting scenarios like person search. On average, the target disappears 38.8 times per sequence, with a maximum of 90 disappearances. Most disappearance intervals range from 0 to 6 seconds, indicating that the target is often occluded but can be re-detected quickly.
During target absences, the sensor’s FoV typically includes 10.6 distractors on average, with some sequences containing more than 20. This highlights the challenge of target re-identification in crowded environments. Additionally, the median target bounding box area is approximately 104.8 $\text{pixel}^{2}$, with a variance of around 164.2 $\text{pixel}^{2}$ per sequence. This variance arises from the target’s free movement, which introduces challenges in tracking and re-identification due to frequent scale changes.
IV Evaluation of TPT Algorithms
In this section, we first present the evaluation protocol and metrics from the perspectives of video object tracking and robotics applications (Sec. IV-A). Next, we introduce baselines that are capable of addressing TPT problems (Sec. IV-B). We then analyze and discuss the experimental results on the TPT-Bench (Sec. IV-C and Sec. IV-D). Additionally, we evaluate trackers on different datasets to demonstrate the TPT-Bench is more challenging with the objective of long-term tracking. (Sec. IV-E). Finally, we discuss potential directions based on the experimental results (Sec. IV-F).
IV-A Evaluation Protocol and Metrics
This dataset can be divided into training and testing subsets; however, in this study, we use all 35 sequences to assess tracking performance, providing a large-scale evaluation of the trackers. Using the initial template, we evaluate tracking performance over sequential frames. Since we target the evaluation of long-term tracking ability, we adopt widely-used metrics in long-term object tracking[16]: tracking precision (TP), tracking recall (TR) and F-score (F).
For each frame at timestep $t$, we have $G_{\_}t$ as the ground-truth target bounding box, $A_{\_}t(\tau_{\_}{\theta})$ as the estimated bounding box by the tracker, $\theta_{\_}t$ the prediction confidence and $\tau_{\_}{\theta}$ a classification threshold. The ground truth is empty if the target person disappears, i.e., $G_{\_}t=\emptyset$. Similarly, the estimation is empty, i.e., $A_{\_}t(\tau_{\_}{\theta})=\emptyset$ if the prediction confidence is below a classification threshold, i.e., $\theta_{\_}t<\tau_{\_}{\theta}$. The agreement between the ground truth and the prediction is quantified using the intersection-over-union (IoU), denoted as $\Omega(A_{\_}t(\tau_{\_}{\theta}),G_{\_}t)$. By definition, this IoU is zero if either $A_{\_}t(\tau_{\_}{\theta})$ or $G_{\_}t$ is empty. A true positive is then determined by whether the overlap $\Omega(A_{\_}t(\tau_{\_}{\theta}),G_{\_}t)$ exceeds a predefined IoU threshold $\tau_{\_}{\Omega}$. Given the two thresholds $(\tau_{\_}{\theta},\tau_{\_}{\Omega})$, TP and TR are defined as:
$\begin{split}&{\textit{TP}}(\tau_{\_}{\theta},\tau_{\_}{\Omega})=|\{t:\Omega(A_{\_}t(\tau_{\_}{\theta}),G_{\_}t)\geq\tau_{\_}{\Omega}\cap A_{\_}t(\tau_{\_}{\theta})\neq\emptyset\}|/N_{\_}p\\ &{\textit{TR}}(\tau_{\_}{\theta},\tau_{\_}{\Omega})=|\{t:\Omega(A_{\_}t(\tau_{\_}{\theta}),G_{\_}t)\geq\tau_{\_}{\Omega}\cap G_{\_}t\neq\emptyset\}|/N_{\_}g\end{split}$ (1)
where ||$\cdot$|| ⋅ | is the cardinality, $N_{\_}g$ is the number of frames with $G_{\_}t\neq\emptyset$, and $N_{\_}p$ is the number of frames with target prediction, i.e., $A_{\_}t(\tau_{\_}{\theta})\neq\emptyset$. To quantify the accuracy of target absence prediction and target ReID capabilities, [16] re-formulates the above equations by integrating over $\tau_{\_}{\Omega}$, resulting in ${\textit{TP}}(\tau_{\_}{\theta})$ and ${\textit{TR}}(\tau_{\_}{\theta})$. Then F is defined as:
$\textit{F}(\tau_{\_}{\theta})=2{\textit{TP}}(\tau_{\_}{\theta}){\textit{TR}}(\tau_{\_}{\theta})/({\textit{TP}}(\tau_{\_}{\theta})+{\textit{TR}}(\tau_{\_}{\theta}))$ (2)
Additionally, we include max recall at 100% precision (MR), a crucial metric[32] for the robotics community often overlooked in existing object tracking benchmarks. This metric evaluates the tracking performance with zero false positives, as false positives are particularly detrimental for robotics systems, where even a false positive could mistakenly lead the robot to follow the wrong person. From the detection literature, the definition of a true positive is tricky since it depends on the IoU threshold $\tau_{\_}{\Omega}$. To solve this, for example, [33] averages performance over various thresholds. Similarly, we average MR over various IoU thresholds. As consequence, average max recall at 100% precision (AMR) is defined as:
$\textit{AMR}=\frac{\sum_{\_}{\tau_{\_}{\Omega}}^{\{\tau_{\_}{\Omega}\}}\max\{\textit{TR}(\tau_{\_}{\theta},\tau_{\_}{\Omega})|\textit{TP}(\tau_{\_}{\theta},\tau_{\_}{\Omega})=1,\tau_{\_}{\theta}\ \in[0,1]\}}{|\{\tau_{\_}{\Omega}\}|}$ (3)
Moreover, we adopt the short-term tracking metric[33], average overlap (AO) for evaluating the tracking accuracy over all frames that the target person is visible, i.e., TR at $\tau_{\_}{\theta}=0$.
IV-B Baselines
For SOT, we build upon the baseline comparisons in [34] and include a range of trackers to comprehensively evaluate long-term tracking performance. Our selection encompasses short-term trackers with online target representation, such as DiMP [35], SiameseRPN++ [36], and TAMOs [37]. Additionally, we evaluate state-of-the-art short-term transformer-based trackers, including ToMP (ResNet101) [38], STARK [39], and MixFormer variants (ConvMAE [40] and CVT [41]). For specialized long-term tracking, we incorporate LTMU [42], which integrates a global tracker (GlobalTrack [43]) with a local tracker (DiMP), and Siam-RCNN [44], which leverages dynamic programming to model historical target paths. We also evaluate KeepTrack [45], a distractor-aware tracker designed for crowded scenarios, to assess its suitability for TPT tasks.
For “MOT + Target-ReID,” we adopt ByteTrack [46] combined with three contemporary Target-ReID methods: CARPE-ID [5], RPF-ReID [4], and RPF-ReID with online continual learning (OCL) [4]. CARPE-ID employs an adaptive target feature updating strategy based on exponential moving averages, while RPF-ReID trains an online ridge-regression classifier using target and non-target features. RPF-ReID with OCL further enhances performance by fine-tuning the feature extractor with online-identified samples (including both target and non-target images), thereby adapting to long-term appearance changes. To investigate the impact of person-ReID capability on TPT performance, we utilize two feature extractors: ResNet18 (R18) pre-trained on MOT20 [11] and KPR [47], a transformer-based partial-ReID method pre-trained on the OCC-Duke dataset [48]. The latter is particularly well-suited for occlusion-rich, robot-egocentric scenarios. Notably, because KPR’s training process relies on human joint annotations—which are challenging to obtain during OCL—we do not apply OCL to this feature extractor.
Finally, we implement a “Detection + ReID” baseline as a conventional pipeline. We use a fixed-size queue to store target features and identify the target person by selecting the detected candidate (by YOLOX[49]) with the highest feature similarity. This approach mirrors standard practices in ReID-driven MOT systems and serves as a foundational comparison point for more advanced methods. As with the above baselines, we adopted two feature extractors (R18 and KPR) for evaluation. Method flows of the above baselines are provided in the supplementary materials.
Method AO F-Score AMR FPS
DiMP[35] 13.11 12.08 0.00 53.9
MixFormer-CVT[40] 14.57 15.19 0.34 35.9
SiameseRPN++[36] 11.78 15.19 2.07 55.4
MixFormer-ConvMAE[41] 21.43 20.72 0.18 18.9
TAMOs[37] 22.28 23.27 0.03 12.9
ToMP[38] 31.37 32.55 0.00 39.1
LTMU[42] 40.60 40.80 3.49 13.1
KeepTrack[45] 37.44 40.90 5.50 28.6
STARK[39] 43.12 43.94 1.37 36.4
Siam-RCNN[44] 46.10 45.03 9.62 3.7
Detection w/ R18[11] 40.92 38.20 3.74 20.7
Detection w/ KPR[47] 55.86 53.22 7.17 10.2
ByteTrack[46]+Target-ReID
CARPE-ID w/ R18 22.04 21.35 0.76 21.2
CARPE-ID[5] w/ KPR 53.61 52.65 11.65 12.0
RPF-ReID w/ R18 50.19 48.38 5.18 20.3
RPF-ReID[4] w/ KPR 65.31 63.26 16.16 11.5
RPF-ReID+OCL[4] w/ R18 57.67 55.71 11.29 18.5
TABLE II: Tracker performance comparison on the TPT-Bench. Off-the-shelf, all trackers perform poorly, demonstrating the new challenges of TPT-Bench. Higher performance from tracking by the type of “ByteTrack[46]+Target-ReID.” The best result is indicated in bold, while the second-best result is underscored.
Method Chen’s dataset[8] LaSOT[14] TPT-Bench
F-Score AMR F-Score AMR F-Score AMR
STARK[39] 97.8 54.0 74.3 32.0 43.9 1.4
Siam-RCNN[44] 94.3 65.9 69.0 30.2 45.0 9.6
Detection w/ KPR[47] 94.4 37.4 58.5 22.9 53.2 7.2
RPF-ReID[4] w/ KPR 99.1 55.5 70.3 37.1 63.3 16.2
RPF-ReID+OCL[4] w/ R18 94.4 47.9 65.6 25.8 55.7 11.3
TABLE III: Tracker performance on different datasets. Compared to existing benchmarks, the trackers on the TPT-Bench exhibit significantly lower performance in terms of F-Score and AMR, highlighting more challenging long-term tracking scenarios.
IV-C Performance of SOT trackers
Baseline results on the TPT-Bench are presented in Table II, and several key observations are worth highlighting. First, long-term trackers demonstrate superior performance, with high F-Scores for LTMU (40.80), KeepTrack (40.90), and Siam-RCNN (45.03). In contrast, most short-term trackers struggle on the TPT-Bench, achieving lower F-Scores, such as DiMP (12.08), MixFormer-ConvMAE (20.72), TAMOs (23.27), and ToMP (32.55). This result is expected as long-term trackers are designed with a focus on re-detection. An exception among short-term trackers is STARK, which achieves the highest F-Score at 43.94, owing to its second training stage that teaches the model to classify the presence of the object.
Second, while STARK achieves a high F-Score, it reports a relatively low AMR of 1.37. In contrast, long-term trackers exhibit higher AMRs, such as LTMU (3.49), KeepTrack (5.50), and Siam-RCNN (9.62). This suggests that a well-designed re-detection mechanism is crucial for the identification and tracking of the target person—an essential capability in robotics, where incorrect person identification can lead to undesired behavior.
Finally, we observe that SOT methods generally underperform on the TPT-Bench when compared to the top-performing methods from other types that leverage prior knowledge about humans. For instance, while Siam-RCNN achieves a higher AMR than “Detection w/ KPR” (+2.45), its F-Score is lower by -8.19. Additionally, its F-Score and AMR are both lower than “CARPE-ID w/ KPR” by -7.62 and -2.03, “RPF-ReID w/ KPR” by -18.23 and -6.54, and “RPF-ReID+OCL w/ R18” by -10.68 and -1.67. This is because SOT methods typically focus on general object tracking and do not consider the use of human-related prior knowledge. Other types of methods incorporate prior knowledge about humans, such as people detection (spatial cues) and person ReID features (appearance cues), which can be more beneficial for TPT.
IV-D Performance of Target-ReID Methods
Target-ReID refers to the methodology used to construct the appearance model of a target person. These appearance models usually leverage pre-trained ReID features. From Table II, we can observe that each method equipped with KPR performs better than that with ResNet18. For example, “Detection w/ KPR” achieves 53.22 F-Score and 7.17 AMR, larger than “Detection w/ R18” by +15.02 F-Score and +3.43 AMR. This is a straightforward verification of the positive impact of ReID ability on TPT since the target person is directly identified based on a feature similarity matching. Furthermore, other methods also benefit from stronger ReID features, e.g., “RPF-ReID w/ KPR” achieves the best result with 63.26 F-Score and 16.16 AMR, larger than “RPF-ReID w/ R18” by +14.88 F-Score and +10.98 AMR.
Leveraging the same feature extractor, RPF-ReID outperforms CARPE-ID. For instance, when equipped with KPR, it surpasses CARPE-ID by +10.61 F-Score and +4.51 AMR; with R18, it exceeds CARPE-ID by +27.03 F-Score and +4.42 AMR. This improvement can be attributed to Ridge Regression, which optimizes the classification boundary by leveraging both target and non-target ReID features collected online. In contrast, CARPE-ID merely aggregates target ReID features in a decaying manner within the original ReID feature space, failing to enhance discrimination further.
Besides learning a classifier online, “RPF-ReID+OCL w/ R18” also fine-tunes the feature extractor using one of the replay-based OCL techniques, Reservoir[50]. This method consolidates the memory by randomly deleting samples based on a probabilistic distribution related to observation times. It then replays training samples from the memory to fine-tune the feature extractor. As a result, it achieves 55.71 F-Score and 11.29 AMR, surpassing “RPF-ReID w/ R18” by +7.33 F-Score and +6.11 AMR.
IV-E Performance on Different Datasets
To highlight the more challenging long-term tracking scenarios in TPT-Bench, we further evaluate several high-performing trackers on other datasets that also involve tracking a person over an extended period (LaSOT[14] and Chen’s dataset[8]). The results are shown in Table III. It is evident that TPT-Bench poses greater long-term tracking challenges compared to the other two datasets, as the best performance achieved by a tracker is only 63.3 F-Score and 16.2 AMR. These values are significantly lower than those in LaSOT, with a drop of -11.0 F-Score and -20.9 AMR, and in Chen’s dataset, with a drop of -35.8 F-Score and -49.7 AMR.
Furthermore, we observe that the best baseline evaluated in TPT-Bench, namely “RPF-ReID w/ KPR,” performs well in the other two person-tracking datasets as well. According to the F-Score metric, it achieves an impressive 99.1 F-Score in Chen’s dataset and a second-best 70.3 F-Score in LaSOT. This is only -4.0 F-Score lower than the best-performing tracker, STARK[39], which is pre-trained on LaSOT. This demonstrates that leveraging prior knowledge of human appearance can significantly enhance the generalizability of TPT in diverse tracking scenarios.
IV-F Discussion
The above experiments suggest “RPF-ReID w/ KPR” to be the most competitive method. We focus on this method for additional analysis, suggesting improvements to long-term egocentric TPT. From experiments, we identify three key takeaways for the potential improvement:
  1. 1.
    Designing better TPT-driven memory consolidation. A failure case shown in Fig. 4 (a) illustrates a scenario where the method fails to re-identify the target from the front-view (third subfigure) until it reappears from the back-view (fourth subfigure). This failure under pose changes may stem from the classifier being trained primarily on abundant back-view observations of the target before it is lost (second subfigure). However, the front-view observation is present in earlier frames (first subfigure). A similar case is shown in Fig. 4 (c), where the classifier fails to re-identify the target after slight clothing changes, even though the reappeared appearance had been observed in previous frames. This highlights the need for improved memory consolidation methods that capture diverse historical samples to better support target re-identification under appearance variation.
  2. 2.
    Improving feature extractor. Although there are no changes in pose or clothing in Fig. 4 (b), lighting variations cause a significant drop in the estimated confidence when the target reappears, indicating insufficient lighting robustness in the pre-trained ReID model. Additionally, in the third subfigure of Fig. 4, where the target is absent, a distractor obtains a high score of 0.57, comparable to that of the true target. To avoid false positives, the classifier must produce high-confidence estimates for the target to reliably distinguish it from distractors. These challenges can be attributed to the non-discriminative representation of the pre-trained ReID features. To improve target discrimination with high confidence, integrating feature uncertainty learning[51] into our previously evaluated state-of-the-art partial ReID method, KPR[47], could be an option to mitigate the impact of noisy training samples. Furthermore, our dataset offers a valuable resource for training more discriminative ReID features.
  3. 3.
    Stablizing classifier training. In Fig. 4 (d), despite minimal visual changes, performance suddenly degrades, with estimated confidence dropping from 0.95 to 0.41. This illustrates the stability-plasticity dilemma[52], where the classifier struggles to maintain stable representations while incorporating new samples. While many methods[52] address this challenge, they have been primarily evaluated on simplified datasets. Our dataset presents a realistic scenario, and adapting existing solutions to the TPT task remains an open question.
V Conclusion
Appendix
V-A Platform, Annotation and Labels
Based on the designed platform, we collect multi-sensor data. Below, we describe the characteristics of each sensor:
@@ V-A Platform, Annotation and Labels
  • Odometry / Actions: Similar to SCAND [29] and MuSoHu [30], we record visual-inertial odometry provided by the ZED 2 camera. This information can be decomposed into linear and angular velocities as the training data for learning-based navigation methods [30].
In addition to the above multi-sensor data, our dataset includes annotations for the target person in panoramic images with 2D bounding boxes. In the following, we will discuss the details of the annotation procedure, bounding box parameterization and additional metadata.
The annotation is performed by all involved students, who are specially instructed for the task. Considering the target’s irregular motion and complex backgrounds, we annotate frames at 15 Hz for accuracy. We adopt an annotation standard similar to LaSOT [14] that if the target person appears in a frame, a labeler should ensure the bounding box is the tightest up-right one to fit any visible part of the target person; otherwise, the labeler gives an absent label, indicating the target person is under full occlusions or invisible.
@@
Next, the annotator reviews each frame, refining any imperfectly estimated bounding boxes if the target is present. If the target does not exist visually, an “absent” label is applied. In cases where the target exists but is misclassified as absent by the Target-ReID method, we first identify the target by selecting its ID from the tracked bounding boxes. Then, we either fine-tune the estimated bounding box or manually draw a new one if no bounding box for the target is available. Finally, we upsample the annotations from 15 Hz to 30 Hz using linear interpolation. For each frame, we record the following metadata (most of them are followed by JRDB [17]):
@@
  • is_behind_glass: A binary indicator that denotes whether the target is behind glass. This annotation is particularly useful in indoor environments with glass walls or transparent obstacles, where depth sensors may misinterpret the target’s distance
  • box: A 2D bounding box defined by the location of the upper-left corner, $(u,v)$, and the size of the bounding box, $(w,h)$.

### v2

@@ 3.1.1 Platform
To facilitate the collection of multi-sensor and egocentric data under robot kinematic constraints for non-robotics experts, we have developed a data collection platform illustrated in Figure 2. This platform consists of a pushing cart equipped with a panoramic camera, a stereo camera, and a 3D LiDAR, operating under differential wheel constraints. The sensors are mounted at typical heights for mobile robots: the panoramic camera is placed approximately 72 cm above the ground, the 3D LiDAR at about 54 cm, and the stereo camera at approximately 45 cm with an elevation angle of around 30°. The extrinsic parameters of these sensors are well-calibrated and provided. Based on the designed platform, we collect multi-sensor data. Below, we describe the characteristics of each sensor:
@@
  • Odometry / Actions: Similar to SCAND (Karnan et al., 2022) and MuSoHu (Nguyen et al., 2023), we record visual-inertial odometry provided by the ZED 2 camera. This information can be decomposed into linear and angular velocities as the training data for learning-based navigation methods (Nguyen et al., 2023).
Although our dataset includes multi-sensor data, we have annotated only the images from the panoramic camera to assess TPT performance using vision-based trackers. We will release all data in ROSBAG format, enabling future research on multi-modal TPT combining visual and depth modalities, natural following through imitation learning with multiple modalities and odometry, and more. Our data collection approach mirrors a human follower’s intention without delay, and all sensors are mounted similarly to the locations of most mobile robots. As such, the collected data can be directly used to train end-to-end networks for performing socially aware companionship, as demonstrated by recent advancements in learning-based social navigation (Karnan et al., 2022; Nguyen et al., 2023).
Figure 2: The platform and sensor suite. A panoramic camera, depth camera, and 3-D LiDAR are fixed on a push-cart in positions that mirror the typical sensor layout of most mobile robots.
Figure 3: Third-view examples of dataset collection procedure. The dataset is demonstrated by a human pushing a sensor-equipped pushcart and following a target person in crowded environments indoors and outdoors. Blue represents the follower, and red represents the target person.
3.1.2 Scenarios
To capture these long-term tracking challenges, we conducted person-following tasks across various indoor and outdoor locations during peak times, such as after classes, during lunch, and at dinner hours. The environments included canteens, laboratories, offices, university building halls, corridors, classrooms, supermarkets, roadside markets, crosswalks, sidewalks, metro stations, and more. Nine students (three females and six males) were recruited as target humans, navigating through these crowded environments and encountering issues like motion blur, varying lighting conditions, changing scales, and alterations in pose and clothing. These students also took on the role of followers, operating the sensor-equipped cart while adhering to social norms, ensuring they did not interrupt the leader or other pedestrians. Additionally, we occasionally carried out person-search tasks (Goldhoorn et al., 2014), which are essential for HRI, by asking the follower to fulfill the leader’s requirements first and then searching for the leader to resume the person-following behavior. Some third-view and robot-egocentric visual examples are shown in Figure 3 and Figure 1, respectively.
As a result, we compiled a large-scale, long-term, multi-modal, and diverse egocentric dataset consisting of 35 sequences captured in different locations, with multiple recordings taken at various times (e.g., daytime, sunset, and night). Some multi-modal examples are shown in Figure 4.
3.2 Annotation and Labels
Our dataset includes annotations for the target person in panoramic images with 2D bounding boxes. In the following, we will discuss the details of the annotation procedure, bounding box parameterization and additional metadata.
The annotation is performed by all involved students, who are specially instructed for the task. Considering the target’s irregular motion and complex backgrounds, we annotate frames at 15 Hz for accuracy. We adopt an annotation standard similar to LaSOT (Fan et al., 2021) that if the target person appears in a frame, a labeler should ensure the bounding box is the tightest up-right one to fit any visible part of the target person; otherwise, the labeler gives an absent label, indicating the target person is under full occlusions or invisible.
@@ 3.2 Annotation and Labels
Next, the annotator reviews each frame, refining any imperfectly estimated bounding boxes if the target is present. If the target does not exist visually, an “absent” label is applied. In cases where the target exists but is misclassified as absent by the Target-ReID method, we first identify the target by selecting its ID from the tracked bounding boxes. Then, we either fine-tune the estimated bounding box or manually draw a new one if no bounding box for the target is available. Finally, we upsample the annotations from 15 Hz to 30 Hz using linear interpolation. For each frame, we record the following metadata (most of them are followed by JRDB (Martin-Martin et al., 2021)):
@@
  • is_behind_glass: An indicator that specifies whether the target is behind glass, where $-1$ denotes the absence of the target, $1$ indicates the target is behind the glass, and 00 means the target is not behind it. This annotation is especially useful in indoor environments with glass walls or transparent obstacles, as depth sensors may incorrectly estimate the target’s distance.
  • bbox: A 2D bounding box is represented as $[u,v,w,h]$, where $(u,v)$ denotes the coordinates of the upper-left corner, and $(w,h)$ represents the width and height of the box. If the target is absent, the bounding box is denoted as $[0,0,0,0]$.